OpenStudy (anonymous):

the circle given by x^2+y^2+4x-10=0 can be written in standard form like this: (x-h)^2+y^2=14 what is the value of h

OpenStudy (nenadmatematika):

h=-2

OpenStudy (anonymous):

h = -2

OpenStudy (anonymous):

(x+2)^2+y^2=10+4 (x+2)^2+y^2=14 so the value of h is -2 is look like +2 but you always negate the value in bracket thus why is h=-2